OpenStudy (anonymous):

A landscaper is designing a flower garden in the shape of a trapezoid. She wants the length of the shorter base to be 3 yards greater than the height, and the length of the longer base to be 7 yards greater than the height. She wants the area to be 225 square yards. The situation is modeled by the equation 1/2h(b1+b2), where h is height in yards, and b1 and b2 are the length of the bases in yards. Complete the square to find the height that will give the desired area.

OpenStudy (anonymous):

@Hero can you help me out again? I'm stuck on this problem... I don't know how to set up the equation

hero (hero):

Okay, basically, the best way to approach this is to start with the formula \[A_{\text{trap}} = \frac{(b_1 + b_2)h}{2}\]

OpenStudy (anonymous):

okay

hero (hero):

Then write \(b_1\) and \(b_2\) in terms of the height. Try doing so.

OpenStudy (anonymous):

are they x+3 and x+7?

hero (hero):

Close: \(b_1 = h + 3\) \(b_2 = h + 7\)

OpenStudy (anonymous):

oooh okay. So now put those into the equation, correct?

hero (hero):

Yes

OpenStudy (anonymous):

\[A=\frac{ (h+3+h+7)h }{ 2 }\]

hero (hero):

Looks good so far

OpenStudy (anonymous):

\[A=\frac{ 2h+10)h }{ 2 }\]

OpenStudy (anonymous):

oops i forgot the parenthesis

hero (hero):

You already know the area as well so we can insert that as well.

OpenStudy (anonymous):

oh right! okay \[225=\frac{ (2h+10)h }{ 2 }\]

hero (hero):

Okay, now try solving for h

OpenStudy (anonymous):

so i distribute the h first, right?

hero (hero):

You can do that, but first I would recommend getting rid of that denominator. Multiply both sides by two.

OpenStudy (anonymous):

okay \[450=(2h+10)h\]

OpenStudy (anonymous):

now distribute the h, correct?

OpenStudy (anonymous):

\[450=2h ^{2} +10h\]

hero (hero):

That looks good so far. What can you do next? Do you know?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

factor maybe?

hero (hero):

Well, it would be a good idea to make sure what you have is in the absolute simplest form first before trying to factor anything.

hero (hero):

We can divide both sides by a number to reduce the equation, correct?

OpenStudy (anonymous):

oh! 2?

hero (hero):

Yup

OpenStudy (anonymous):

\[225=x ^{2}+5\]

OpenStudy (anonymous):

i meant 5x

OpenStudy (anonymous):

or h hahaha

hero (hero):

You shouldn't change the \(h's\) to \(x's\)

hero (hero):

Yes try to keep track of the variables and other things

OpenStudy (anonymous):

yeah, sorry I'm just used to x, my mistake

hero (hero):

It's okay. It happens.

OpenStudy (anonymous):

so now we factor?

hero (hero):

So we have \(225 = h^2 + 5h\)

hero (hero):

Right but we have to put it in an appropriate form first before factoring. Or perhaps at this point they want you to complete the square

OpenStudy (anonymous):

225=(x+5)(x+0)

OpenStudy (anonymous):

oh...so how do we do that?

hero (hero):

Basically, when you have a quadratic equation of the form \(x^2 + bx = c\) You have to add\(\left(\dfrac{b}{2}\right)^2\) to both sides. In this case, b = 5 and c = 225

OpenStudy (anonymous):

okay....

OpenStudy (anonymous):

Sorry, you kind of lost me

hero (hero):

To complete the square, you simply add \(\left(\dfrac{b}{2}\right)^2\) to both sides. In this case, b = 5 to add \(\left(\dfrac{5}{2}\right)^2\) to both sides.

OpenStudy (anonymous):

ooooh okay

OpenStudy (anonymous):

so now the equation would be

OpenStudy (anonymous):

so now the equation would be\[(\frac{ 5 }{ 2 })^{2}\times225=(\frac{ 5 }{ 2 })x^{2}+5x\]

hero (hero):

you'rere supposed to \(\color\red{\text{ADD}}\) \(\left(\dfrac{5}{2}\right)^2\) to both sides.

hero (hero):

not multiply